OpenStudy (anonymous):

for the function \(y=2-4e^-(x-1)^2\) find the equation of the tangent line to the function y at =2

OpenStudy (anonymous):

\[y=2-4e ^{-(x-1)^{2}}\]

OpenStudy (campbell_st):

can you find the 1st derivative..?

OpenStudy (anonymous):

\[8(x-1)e ^{-(x-1)^{2}}\]

OpenStudy (campbell_st):

ok... so this is the 1st derivative and also the equation of the slope of the tangent so substitute x = 2.... ( I think you left x = 2) out of the question

OpenStudy (anonymous):

\[8e ^{-1}\]

OpenStudy (anonymous):

does that look right?

OpenStudy (campbell_st):

great so that is the slope now you need a point on the curve substitute x = 2 into the original equation. to get a y value

OpenStudy (anonymous):

3.4715?

OpenStudy (campbell_st):

but you have a slight error in the equation for the slope \[y = 2 - 4e^{-(x -1)^2}\] then \[\frac{dy}{dx} = -8(x -1)e^{-(x -1)^2}\] can you leave the y value for the point as an exact value..?

OpenStudy (campbell_st):

so substituting x = 2 the slope is \[m = -8e^{-1}\]

OpenStudy (anonymous):

but isn't -2x-4=8?

OpenStudy (campbell_st):

now the point substituting x = 2 and getting an exact value its \[y = 2 - 4e^{-1}\] so now you have a point \[(2, 2 - 4e^{-1})\] and a slope... so you can find the equation of the tangent using the point slope formula or slope intercept formula. hope it helps

OpenStudy (campbell_st):

oops you are correct... my mistake on the derivative

OpenStudy (anonymous):

what answer did you get for 2-4e^-1